# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

### v3

@@ Abstract
The rapid scaling of artificial intelligence models has revealed a fundamental tension between model capacity (storage) and inference efficiency (computation). While classical information theory focuses on transmission and storage limits, it lacks a unified physical framework to quantify the thermodynamic costs of generating information from compressed laws versus retrieving it from memory. In this paper, we propose a theoretical framework that treats information processing as an enabling mapping from ontological states to carrier states. We introduce a novel metric, Derivation Entropy, which quantifies the effective work required to compute a target state from a given logical depth. By analyzing the interplay between Shannon entropy (storage) and computational complexity (time/energy), we demonstrate the existence of a critical phase transition point. Below this threshold, memory retrieval is thermodynamically favorable; above it, generative computation becomes the optimal strategy. This ”Energy-Time-Space” conservation law provides a physical explanation for the efficiency of generative models and offers a rigorous mathematical bound for designing next-generation, energy-efficient AI architectures. Our findings suggest that the minimization of Derivation Entropy is a governing principle for the evolution of both biological and artificial intelligence.
@@ I Introduction
Classically, these two aspects have been treated separately. Shannon’s information theory [shannon1948mathematical] rigorously defines the limits of static storage and transmission, quantifying the irreducible randomness of the carrier state. However, Shannon entropy is agnostic to the causal history or the computational effort required to generate that state. Conversely, Algorithmic Information Theory (AIT), particularly through the lens of Kolmogorov complexity [kolmogorov1965three] and Bennett’s concept of Logical Depth [bennett1988logical], focuses on the computational time required to reconstruct a structure from its compressed description. While Logical Depth captures the dynamic value of the ”mapping process,” it lacks a direct unification with the statistical mechanics governing the carrier’s physical existence.
@@
In this paper, we propose the ”Information Physics of Intelligence,” a theoretical framework that unifies Logical Depth and Shannon Entropy under rigorous thermodynamic constraints. We define Derivation Entropy ($H_{derive}$), which quantifies the total physical cost of the ontological-to-carrier mapping. By integrating the thermodynamic cost of storage with the algorithmic complexity of retrieval, we derive the Energy-Time-Space Triality Bound. This bound reveals that minimizing storage space (lowering Shannon entropy) inevitably necessitates a compensatory increase in computational time (increasing Logical Depth) or energy flux, governed by Landauer’s limit.
@@ II-A The Underlying Logical System $\mathcal{L}$
  • Expressive adequacy: By the Immerman–Vardi theorem, FO(LFP) captures exactly the class of polynomial-time computable properties over ordered finite structures. This makes $\mathcal{L}$ sufficiently expressive to model the state evolution of a wide range of real-world information systems under finite-horizon autoregressive decoding.
@@ II-C Information Cost of Energy and Time
Summing the resource inequality over all $N$ transitions yields
$E\cdot\tau\;\geq\;\sum_{i=1}^{N}E_{i}\cdot\tau_{i}\;\geq\;N\kappa\;\geq\;\kappa I,$
where $E$ and $\tau$ denote, respectively, the total energy resource and the total implementation time available to the physical process. This establishes the claimed lower bound (5). ∎
@@
In a thermal environment at temperature $T$, Landauer’s principle[landauer1961irreversibility] states that any logically irreversible operation that erases or thermodynamically commits one bit of information leads, in the quasistatic limit, to a minimal heat dissipation of at least $k_{B}T\ln 2$. In the setting of Theorem 1, implementing an ideal information instance that reliably encodes $I=\log_{2}|S_{O}(O,T_{o})|$ bits requires at least $I$ effective information-bearing operations, each of which is thermodynamically equivalent to committing one bit of semantic information to a physical carrier state.
@@ II-F Derivation Depth
With the atomic basis and the semantic inference structure in place, we can now define the complexity of any derived state in terms of its distance from the atomic layer along the causal/inference graph.
@@
where $\mathrm{depth}_{t}(x\mid y)$ denotes Bennett’s logical depth of string $x$ at significance level $t$, conditional on $y$, with respect to the universal machine $U$.
@@ III Derivation Depth as an Information Metric
We now establish that derivation depth is an information-theoretic quantity dual to Shannon entropy, bridging the gap between logical reasoning and information processing.
@@ III-A Derivation Entropy and Information Content
    Classical thermodynamics: Each logically irreversible step erases $\geq 1$ bit. By Landauer’s principle, $\lambda=\ln 2$ nats.
@@
    Quantum mechanics: By the Margolus–Levitin bound, $\lambda=\pi/2$ in units of $\hbar$.
@@
where $K(\cdot\mid\cdot)$ denotes conditional Kolmogorov complexity with respect to a fixed universal Turing machine and a computable encoding scheme $\langle\cdot\rangle$.
@@
(e.g., Datalog, SPARQL, and template-based natural-language queries), conditions (17)–(18) are asymptotically equivalent to
@@
Let $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$. By Lemma 1 and Lemma 2, there exist constants $c_{1},c_{2}>0$ such that
@@
III-B Frequency-weighted storage–computation tradeoff
In this subsection we combine the semantic information measures developed in Section II with the derivation-depth analysis in Section III.A to obtain a quantitative tradeoff between storage and computation under a given query distribution. Intuitively, if a query $q$ will be accessed frequently, it may be advantageous to “precompute and store” part of the information needed to answer $q$; if $q$ is rare, it is better to compute the answer on demand from the base knowledge $S_{O}$. Our main result in this subsection, Theorem 5, formalizes this intuition in terms of semantic information content, derivation entropy, and frequency-weighted amortized cost.
@@ III-B Frequency-weighted storage–computation tradeoff
$f_{c}\;\asymp\;L_{S}\;=\;\log(m_{S}+n_{S}).$ (35)
Queries with $f_{q}\gg f_{c}$ are asymptotically more cost-effective to support via storage (larger $S$); queries with $f_{q}\ll f_{c}$ are asymptotically more cost-effective to answer via on-demand derivation.
@@
Step 0: Information-theoretic preliminaries
By (SA1) and the data-processing inequality,
@@
Step 1: Applying Theorem 4
By (SA2) and (SA3), Theorem 4 applies to $S_{O}\cup S$ and yields
@@
Step 2: Amortized cost
By definition,
@@
Step 3: Extremal strategies
To obtain a clean upper bound in terms of $(H_{q},f_{q},L_{S})$ alone, we compare against two constructive extremes.
@@
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\max\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1),$ (49)
@@
Step 4: Frequency regimes
From the last inequality we obtain two asymptotic regimes:
@@
    If we model the “effective” per-parameter capacity at the scale of GPT-4 ($N\approx 1.7\times 10^{12}$) by $\log N\approx 40$ bits, this suggests that only facts whose derivation requires on the order of tens of bits of computation and that are accessed at least moderately often should be encoded in the weights. Rare or computationally easy facts are better left to retrieval or external tools.
@@
    For a large knowledge graph such as Freebase ($|E|\approx 10^{9}$), this gives $f_{c}\approx 1.05$, meaning that even slightly more-than-once multi-hop patterns can already justify precomputation.
@@
III-C Mutual Information and Submodularity
Theorem 6 (Storage Efficiency via Mutual Information).
The efficiency of storage set $S$ for query distribution $P_{Q}$ is:
$\eta(S)=\frac{I(S;Q)}{|S|}\in[0,1]$ (68)
with equality at 1 if and only if $S$ is a sufficient statistic for $Q$.
Theorem 7 (Submodularity of Mutual Information).
The mutual information $I(S;Q)$ is submodular in $S$:
$I(S_{1}\cup S_{2};Q)+I(S_{1}\cap S_{2};Q)\leq I(S_{1};Q)+I(S_{2};Q)$ (69)
This follows from the submodularity of entropy. For any sets $A,B$:
$H(A\cup B)+H(A\cap B)$ $\leq H(A)+H(B)$ (70)
$\Rightarrow I(A\cup B;Q)+I(A\cap B;Q)$ $\leq I(A;Q)+I(B;Q)$ (71)
by expanding mutual information as I(X;Q)=H(Q)H(Q|X)I(X;Q)=H(Q)-H(Q|X) and using the fact that conditional entropy is supermodular. ∎
Corollary 5 (Greedy Approximation).
A greedy algorithm for selecting storage items achieves a (11/e)(1-1/e) approximation ratio for maximizing $I(S;Q)$ subject to $|S|\leq M_{\max}$.

### v4

# Information Physics of Intelligence: Unifying Derivation Depth and Entropy under Thermodynamic Constraints

@@ Abstract
The rapid scaling of artificial intelligence models has revealed a fundamental tension between model capacity (storage) and inference efficiency (computation). While classical information theory focuses on transmission and storage limits, it lacks a unified physical framework to quantify the thermodynamic costs of generating information from compressed laws versus retrieving it from memory. In this paper, we propose a theoretical framework that treats information processing as an enabling mapping from ontological states to carrier states. We introduce a novel metric, Derivation Entropy, which quantifies the effective work required to compute a target state from a given derivation depth. By analyzing the interplay between Shannon entropy (storage) and computational complexity (time/energy), we demonstrate the existence of a critical phase transition point. Below this threshold, memory retrieval is thermodynamically favorable; above it, generative computation becomes the optimal strategy. This ”Energy-Time-Space” conservation law provides a physical explanation for the efficiency of generative models and offers a rigorous mathematical bound for designing next-generation, energy-efficient AI architectures. Our findings suggest that the minimization of Derivation Entropy is a governing principle for the evolution of both biological and artificial intelligence.
@@ I Introduction
Classically, these two aspects have been treated separately. Shannon’s information theory [shannon1948mathematical] rigorously defines the limits of static storage and transmission, quantifying the irreducible randomness of the carrier state. However, Shannon entropy is agnostic to the causal history or the computational effort required to generate that state. Conversely, Algorithmic Information Theory (AIT), particularly through the lens of Kolmogorov complexity [kolmogorov1965three] and Bennett’s concept of Logical Depth [bennett1988logical], focuses on the computational time required to reconstruct a structure from its compressed description. While Derivation Depth captures the dynamic value of the ”mapping process,” it lacks a direct unification with the statistical mechanics governing the carrier’s physical existence.
@@
In this paper, we propose the ”Information Physics of Intelligence,” a theoretical framework that unifies Derivation Depth and Shannon Entropy under rigorous thermodynamic constraints. We define Derivation Entropy ($H_{derive}$), which quantifies the total physical cost of the ontological-to-carrier mapping. By integrating the thermodynamic cost of storage with the algorithmic complexity of retrieval, we derive the Energy-Time-Space Triality Bound. This bound reveals that minimizing storage space (lowering Shannon entropy) inevitably necessitates a compensatory increase in computational time (increasing Derivation Depth) or energy flux, governed by Landauer’s limit.
@@ II-A The Underlying Logical System $\mathcal{L}$
  • Expressive adequacy: By the Immerman–Vardi theorem [immerman1986relational, vardi1982complexity], FO(LFP) captures exactly the class of polynomial-time computable properties over ordered finite structures. This makes $\mathcal{L}$ sufficiently expressive to model the state evolution of a wide range of real-world information systems under finite-horizon autoregressive decoding.
@@ II-C Information Cost of Energy and Time
We analyze two implementation paradigms:
Case 1: Sequential execution. The total time is $\tau=\sum_{i=1}^{N}\tau_{i}$, and energy can be reused across steps, so the effective energy is $E=\max_{i}E_{i}$. By Assumption 1, $E\cdot\tau_{i}\geq\kappa$ for each $i$, hence
$E\cdot\tau=E\cdot\sum_{i=1}^{N}\tau_{i}\geq N\kappa\geq\kappa I.$
Case 2: Parallel execution. The total time is $\tau=\max_{i}\tau_{i}$, and energy must be summed, so $E=\sum_{i=1}^{N}E_{i}$. By Assumption 1, $E_{i}\cdot\tau\geq\kappa$ for each $i$, hence
$E\cdot\tau=\tau\cdot\sum_{i=1}^{N}E_{i}\geq N\kappa\geq\kappa I.$
In both cases, $E\cdot\tau\geq\kappa I$. ∎
@@
In a thermal environment at temperature $T$, Landauer’s principle[landauer1961irreversibility, berut2012experimental] states that any logically irreversible operation that erases or thermodynamically commits one bit of information leads, in the quasistatic limit, to a minimal heat dissipation of at least $k_{B}T\ln 2$. In the setting of Theorem 1, implementing an ideal information instance that reliably encodes $I=\log_{2}|S_{O}(O,T_{o})|$ bits requires at least $I$ effective information-bearing operations, each of which is thermodynamically equivalent to committing one bit of semantic information to a physical carrier state.
@@ II-F Derivation Depth
With the atomic basis and the semantic inference structure in place, we can now define the complexity of any derived state in terms of its distance from the atomic layer along the causal/inference graph [Buss1987Boolean][Krajicek1995Bounded].
@@
where $\mathrm{depth}_{t}(x\mid y)$ denotes Bennett’s logical depth [bennett1988logical] of string $x$ at significance level $t$, conditional on $y$, with respect to the universal machine $U$.
@@ III Derivation Depth as an Information Metric
We now establish that derivation depth is an information-theoretic quantity dual to Shannon entropy [shannon1948mathematical], bridging the gap between logical reasoning and information processing.
@@ III-A Derivation Entropy and Information Content
    Classical thermodynamics: Each logically irreversible step erases $\geq 1$ bit. By Landauer’s principle, $\lambda=\ln 2$ nats [landauer1961irreversibility].
@@
    Quantum mechanics: By the Margolus–Levitin bound, $\lambda=\pi/2$ in units of $\hbar$ [Margolus1998maximum].
@@
Throughout the remainder of this paper, unless otherwise specified, we adopt the classical thermodynamic regime with $\lambda=\ln 2$ nats, so that derivation entropy is measured in the same units (nats) as Shannon entropy.
@@
where $K(\cdot\mid\cdot)$ denotes conditional Kolmogorov complexity with respect to a fixed universal Turing machine and a computable encoding scheme $\langle\cdot\rangle$ [li2008introduction].
@@
(e.g., Datalog [ceri1989you], SPARQL [perez2009semantics]), and template-based natural-language queries), conditions (17)–(18) are asymptotically equivalent to
@@
Let $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$. By Lemma 1 and 2, there exist constants $c_{1}$, $c_{2}>0$, or more precisely, $c_{1}$ and $c_{2}$ depend only on the universal Turing machine, proof system $\mathcal{P}$, and encoding scheme, but not on the polynomial degree in the assumption $n={\rm poly}(m)$, such that
@@
III-B Frequency-weighted Storage–Computation Tradeoff for a Single Query
In this subsection we combine the semantic information metrics to obtain a quantitative tradeoff between storage and computation under a given query distribution.
@@ III-B Frequency-weighted Storage–Computation Tradeoff for a Single Query
$f_{c}\asymp L_{S}=\log(m_{S}+n_{S}),$ (35)
where, for the baseline case $S=\emptyset$, we have $f_{c}=\log(|\mathrm{Atom}(S_{O})|+D_{d}(q|S_{O}))$; see Remark 4 for the general case.
@@
Step 0: Information-theoretic preliminaries By (SA1) and the data-processing inequality,
@@
Step 1: Applying Theorem 4. By (SA2) and (SA3), Theorem 4 applies to $S_{O}\cup S$ and yields
@@
Step 2: Amortized cost. By definition,
@@
Step 3 Extremal strategies. To obtain a clean upper bound in terms of $(H_{q},f_{q},L_{S})$ alone, we compare against two constructive extremes.
@@
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\min\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1),$ (49)
@@
Step 4: Frequency regimes. From the last inequality we obtain two asymptotic regimes:
@@
Remark 4 (Operational interpretation of critical frequency).
The critical frequency $f_{c}$ admits the following operational interpretation:
  1. 1.
    Baseline case ($S=\emptyset$): When no auxiliary storage is used, the critical frequency is determined solely by the base knowledge:
    $f_{c}^{(0)}:=\log\bigl(|\mathrm{Atom}(S_{O})|+D_{d}(q|S_{O})\bigr).$
    This is the threshold for deciding whether any storage investment for query $q$ is worthwhile.
  2. 2.
    Post-allocation case: For a given storage strategy $\Pi=(S,C,R)$, the effective critical frequency shifts to
    $f_{c}(S):=L_{S}=\log(m_{S}+n_{S}),$
    where $m_{S}=|\mathrm{Atom}(S_{O}\cup S)|$ and $n_{S}=D_{d}(q|S_{O}\cup S)$.
  3. 3.
    Decision rule: The initial decision of whether to allocate storage for $q$ should compare $f_{q}$ against the baseline $f_{c}^{(0)}$, not against $f_{c}(S)$, to avoid circular reasoning. Once $S$ is chosen, $f_{c}(S)$ governs the marginal benefit of additional storage.
@@
In summary, Theorem 5 provides a frequency-weighted amortized cost bound at the level of a single query type, while Remark 6 clarifies how the one-time storage cost $|S|$ should be correctly amortized. This completes the frequency-weighted analysis at the level of a single query type.
III-C System-Wide Storage Allocation under Amortized Cost
We now move from the single-query amortization view to a system-wide analysis over a full query distribution.
@@ III-C System-Wide Storage Allocation under Amortized Cost
    If we model the “effective” per-parameter capacity at the scale of GPT-4 ($N\approx 1.7\times 10^{12}$[openai2023gpt4] by $\log N\approx 40$ bits, this suggests that only facts whose derivation requires on the order of tens of bits of computation and that are accessed at least moderately often should be encoded in the weights. Rare or computationally easy facts are better left to retrieval or external tools.
@@
    For a large knowledge graph such as Freebase ($|E|\approx 10^{9}$[bollacker2008freebase], this gives $f_{c}\approx 1.05$, meaning that even slightly more-than-once multi-hop patterns can already justify precomputation.
@@
III-D Mutual Information and Submodularity
The system-wide cost decomposition in Section III.C shows that, for a query workload $Q\sim P_{Q}$ and a storage set $S$, the expected per-query cost can be written as
$\mathbb{E}[\text{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\!\big[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\big]\\ \;\approx\;\frac{|S|}{N}+H(Q\mid S_{O},S),$ (68)
where $H(\cdot\mid\cdot)$ denotes Shannon conditional entropy. Using the identity
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$ (69)
we see that, under a fixed base knowledge $S_{O}$ and query distribution $P_{Q}$, choosing an effective storage set $S$ amounts to maximizing the conditional mutual information $I(S;Q\mid S_{O})$ under storage constraints. Intuitively, $I(S;Q\mid S_{O})$ measures how much additional structure about the future query workload $Q$ is captured by $S$ beyond what is already encoded in $S_{O}$. In this subsection we study mutual information as a set function of $S$, and show that it is submodular, which enables efficient approximate optimization of storage allocation.
We first formalize an efficiency measure that quantifies the information-theoretic benefit of storage per unit capacity consumed.
Definition 10 (Storage efficiency via conditional mutual information).
Fix a base knowledge $S_{O}$ and a query distribution $P_{Q}$ over a discrete query space $\mathcal{Q}$. For any storage set $S$ (viewed as a random object induced by a storage policy), we define its storage efficiency as
$\eta(S)\;:=\;\frac{I(S;Q\mid S_{O})}{|S|},$ (70)
where $I(S;Q\mid S_{O})$ is the conditional mutual information between $S$ and $Q$ given $S_{O}$, and $|S|$ is the bit-length of an effective encoding of $S$ as in assumption (SA1).
The upper bound $\eta(S)\leq 1$ follows from the data-processing inequality and assumption (SA1), which ensures $|S|\leq H(S)+O(1)$. In particular, $\eta(S)=1$ if and only if $S$ is a sufficient statistic for $Q$ given $S_{O}$ and the encoding of $S$ is Shannon-optimal.
The numerator $I(S;Q\mid S_{O})$ in (70) measures how much the storage $S$ reduces the Shannon uncertainty of the query workload beyond what is already known from $S_{O}$:
$I(S;Q\mid S_{O})\;=\;H(Q\mid S_{O})-H(Q\mid S_{O},S).$ (71)
A larger value of $I(S;Q\mid S_{O})$ means that the residual entropy $H(Q\mid S_{O},S)$—and hence the expected inference burden in (68)—is smaller.
The denominator $|S|$ corresponds to the one-time storage cost of $S$ (in bits). Thus $\eta(S)$ can be interpreted as the information gain per unit storage: a larger $\eta(S)$ indicates that each bit of storage allocated to $S$ carries more predictive information about the query workload $Q$ beyond the base knowledge $S_{O}$.
We can make this interpretation explicit in the idealized limit where $S$ serves as a sufficient statistic for $Q$ relative to $S_{O}$.
Proposition 2 (Efficiency and sufficiency).
Fix $S_{O}$ and a query distribution $P_{Q}$. Suppose $S$ is a sufficient statistic for $Q$ given $S_{O}$ in the sense that
$H(Q\mid S_{O},S)=0.$ (72)
Then
$I(S;Q\mid S_{O})=H(Q\mid S_{O}),$ (73)
and hence the storage efficiency
$\eta(S)\;=\;\frac{H(Q\mid S_{O})}{|S|}$ (74)
attains its maximal possible value under the given encoding of $S$.
By the definition of conditional mutual information we have
$I(S;Q\mid S_{O})\;=\;H(Q\mid S_{O})-H(Q\mid S_{O},S).$ (75)
If $S$ is sufficient for $Q$ given $S_{O}$ in the sense of (72), then $H(Q\mid S_{O},S)=0$, and therefore
$I(S;Q\mid S_{O})\;=\;H(Q\mid S_{O}).$ (76)
Since, for any encoding of $S$, we have the generic upper bound
$I(S;Q\mid S_{O})\;\leq\;H(S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|,$ (77)
it follows that, for a fixed encoding of $S$, the ratio $I(S;Q\mid S_{O})/|S|$ is maximized if and only if $I(S;Q\mid S_{O})=H(Q\mid S_{O})$. Under (72), this yields
$\eta(S)\;=\;\frac{I(S;Q\mid S_{O})}{|S|}\;=\;\frac{H(Q\mid S_{O})}{|S|},$ (78)
which is the maximal efficiency consistent with the chosen encoding and storage budget. This proves the claim. ∎
Proposition 2 shows that, in the ideal case where $S$ captures all the uncertainty of $Q$ beyond $S_{O}$, the storage efficiency $\eta(S)$ reaches its conceptual maximum, with each stored bit contributing as much predictive information as the encoding allows. In realistic systems, $S$ will only capture part of the structure of $P_{Q}$, leading to intermediate values of $\eta(S)$ that quantify the tradeoff between storage capacity and workload-specific informativeness.
From the system-level viewpoint of Section III.C, a natural storage allocation problem is:
Given a storage budget $M_{\max}$ (in bits), select a storage set $S$ so as to minimize the expected per-query cost $\mathbb{E}[\text{Cost}]$ in (68), equivalently to maximize the conditional mutual information $I(S;Q\mid S_{O})$ subject to $|S|\leq M_{\max}$.
To analyze this optimization problem we now show that $I(S;Q\mid S_{O})$, viewed as a set function of $S$, is submodular. This structure underlies the greedy approximation guarantee stated later.
Throughout this subsection we treat $S$ as a finite subset of some ground set $\mathcal{U}$ indexing elementary storage items (e.g., candidate facts, cached answers, or intermediate derivations). For any $S\subseteq\mathcal{U}$, we write $I(S;Q\mid S_{O})$ as a shorthand for $I(X_{S};Q\mid S_{O})$, where $X_{S}$ denotes the random vector of storage variables corresponding to the items in $S$.
Proposition 3 (Submodularity of conditional mutual information).
Fix a base knowledge $S_{O}$ and a query random variable $Q$. For any two storage sets $S_{1},S_{2}\subseteq\mathcal{U}$, the conditional mutual information satisfies
$I(S_{1}\cup S_{2};Q\mid S_{O})\;+\;I(S_{1}\cap S_{2};Q\mid S_{O})\\ \;\leq\;I(S_{1};Q\mid S_{O})\;+\;I(S_{2};Q\mid S_{O}).$ (79)
In particular, the set function
$f(S):=I(S;Q\mid S_{O})$ (80)
is submodular in $S$.
Recall that, for any random variables $X,Q$ and conditioning variable $Z$, the conditional mutual information is
$I(X;Q\mid Z)\;=\;H(Q\mid Z)-H(Q\mid Z,X),$ (81)
where $H(\cdot\mid\cdot)$ denotes conditional entropy.
Let us denote, for convenience,
$A:=S_{1},\qquad B:=S_{2}.$ (82)
Then
$I(A\cup B;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},A\cup B),$ (83)
$I(A\cap B;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},A\cap B),$ (84)
$I(A;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},A),$ (85)
$I(B;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},B).$ (86)
Summing the left-hand side of (79) and using (83)–(84) yields
$I(A\cup B;Q\mid S_{O})+I(A\cap B;Q\mid S_{O})$
$\qquad=2H(Q\mid S_{O})-H(Q\mid S_{O},A\cup B)$
$\qquad-H(Q\mid S_{O},A\cap B).$ (87)
Similarly, summing the right-hand side of (79) and using (85)–(86) gives
$I(A;Q\mid S_{O})+I(B;Q\mid S_{O})$
$\qquad=2H(Q\mid S_{O})-H(Q\mid S_{O},A)-H(Q\mid S_{O},B).$ (88)
Therefore, the inequality (79) is equivalent to
$H(Q\mid S_{O},A\cup B)\;+\;H(Q\mid S_{O},A\cap B)$
$\qquad\;\geq\;H(Q\mid S_{O},A)\;+\;H(Q\mid S_{O},B).$ (89)
That is, we need to show that the mapping
$S\;\mapsto\;H(Q\mid S_{O},S)$ (90)
is supermodular in $S$.
This is a standard property of conditional entropy. Intuitively, as we condition on a larger set of variables, the marginal benefit of conditioning on any additional variable decreases, which is precisely the notion of supermodularity for $H(Q\mid S_{O},S)$ and, via (81), submodularity for $I(S;Q\mid S_{O})$.
More formally, for any $A\subseteq B\subseteq\mathcal{U}$ and any element $x\in\mathcal{U}\setminus B$, the marginal reduction in conditional entropy due to conditioning on $x$ satisfies
$H(Q\mid S_{O},A)-H(Q\mid S_{O},A\cup\{x\})$
$\quad=I(x;Q\mid S_{O},A),$ (91)
and
$H(Q\mid S_{O},B)-H(Q\mid S_{O},B\cup\{x\})$
$\quad=I(x;Q\mid S_{O},B).$ (92)
Since conditioning cannot increase mutual information, we have
$I(x;Q\mid S_{O},B)\;\leq\;I(x;Q\mid S_{O},A),$ (93)
which implies that the marginal entropy reduction $H(Q\mid S_{O},S)-H(Q\mid S_{O},S\cup\{x\})$ is decreasing as $S$ grows. This is exactly the defining property of supermodularity of $H(Q\mid S_{O},S)$ in $S$, which in turn is equivalent to the submodularity [fujishige1978polymatroidal] of $I(S;Q\mid S_{O})$ via (81)–(89). This proves (79) and establishes that $f(S)$ in (80) is submodular. ∎
Proposition 3 shows that the incremental utility of adding a storage item to SS—measured by the gain in conditional mutual information $\Delta f:=I(S\cup\{x\};Q\mid S_{O})-I(S;Q\mid S_{O})$—exhibits diminishing returns as $S$ grows. This property is fundamentally important for storage allocation, because it implies that simple greedy procedures enjoy strong approximation guarantees when maximizing $I(S;Q\mid S_{O})$ under a budget constraint.
We now state a direct consequence of Proposition 3 for the optimization problem induced by the system-level cost decomposition (68)–(69).
Corollary 5 (Greedy approximation for storage selection).
Consider the problem of selecting a storage set $S\subseteq\mathcal{U}$ under a budget constraint
$|S|\;\leq\;M_{\max},$ (94)
so as to maximize the conditional mutual information
$f(S):=I(S;Q\mid S_{O}),$ (95)
where $S_{O}$ and $P_{Q}$ are fixed. Assume that the ground set $\mathcal{U}$ is finite, that $f(\varnothing)=0$, and that $f(S)$ is nonnegative and monotone non-decreasing in $S$. Then the standard greedy algorithm that iteratively adds the storage item with the largest marginal gain in $f(S)$ subject to the remaining budget achieves a (11/e)(1-1/e)-approximation to the optimal value:
$f(S_{\mathrm{greedy}})\;\geq\;(1-1/e)\,f(S^{\star}),$ (96)
where $S^{\star}$ is an optimal solution with $|S^{\star}|\leq M_{\max}$, and $S_{\mathrm{greedy}}$ is the solution returned by the greedy procedure.
By Proposition 3, the function $f(S)=I(S;Q\mid S_{O})$ is submodular in $S$. Nonnegativity and monotonicity follow from basic properties of mutual information:
$0\;\leq\;I(S;Q\mid S_{O})\;\leq\;I(S\cup T;Q\mid S_{O})$ (97)
whenever $S\subseteq S\cup T$. Therefore, $f(S)$ satisfies the standard conditions of the classical result on maximizing a nonnegative monotone submodular function under a cardinality or knapsack constraint. By the well-known theorem of Nemhauser et al. [nemhauser1978analysis], the greedy algorithm that repeatedly adds the element with largest marginal gain in $f(S)$ subject to the remaining budget achieves a (11/e)(1-1/e)-approximation to the optimal value. Applying this general result to the specific choice $f(S)=I(S;Q\mid S_{O})$ yields the claimed inequality. ∎
Corollary 5 provides a principled algorithmic prescription for system-wide storage allocation: under the amortized cost model of Section III.C, one should select storage items greedily in order of their marginal contribution to $I(S;Q\mid S_{O})$ per unit of storage cost, subject to the overall budget $M_{\max}$. Up to the universal (11/e)(1-1/e) factor, this strategy is provably as effective as any other storage allocation policy in terms of reducing the expected inference burden $\mathbb{E}_{q\sim P_{Q}}[H_{\mathrm{derive}}(q\mid S_{O}\cup S)]$ and, equivalently, the Shannon term $H(Q\mid S_{O},S)$ in (68).